### v1

# Full-scale Representation Guided Network for Retinal Vessel Segmentation

Abstract
The U-Net architecture and its variants have remained state-of-the-art (SOTA) for retinal vessel segmentation over the past decade. In this study, we introduce a Full Scale Guided Network (FSG-Net), where the feature representation network with modernized convolution blocks extracts full-scale information and the guided convolution block refines that information. Attention-guided filter is introduced to the guided convolution block under the interpretation that the filter behaves like the unsharp mask filter. Passing full-scale information to the attention block allows for the generation of improved attention maps, which are then passed to the attention-guided filter, resulting in performance enhancement of the segmentation network. The structure preceding the guided convolution block can be replaced by any U-Net variant, which enhances the scalability of the proposed approach. For a fair comparison, we re-implemented recent studies available in public repositories to evaluate their scalability and reproducibility. Our experiments also show that the proposed network demonstrates competitive results compared to current SOTA models on various public datasets. Ablation studies demonstrate that the proposed model is competitive with much smaller parameter sizes. Lastly, by applying the proposed model to facial wrinkle segmentation, we confirmed the potential for scalability to similar tasks in other domains. Our code is available on https://github.com/ZombaSY/FSG-Net-pytorch.
1 Introduction
Convolutional neural networks (CNNs) have seen significant improvements in performance and optimization since the 2010s. The introduction of hardware acceleration using GPUs, ReLU activation function, and residual block [11, 1] has enabled smooth back-propagation in deep neural network architectures. Research in this field has focused on finding a balance between computational efficiency, parameter size, code scalability, and fidelity. Depthwise separable convolution [5] and squeeze-and-excitation [12] have been particularly influential in this regard. The inverted residual block [19] achieved higher fidelity with optimized computational efficiency and a smaller parameter size compared to ResNet.
Figure 1: F1 scores of compared networks on the DRIVE dataset, measured against the validation dataset comprising zero-padded images of resolution 608$\times$608. Among the considered architectures, FSG-Net-T achieved a superior F1 score compared to competitive models while maintaining a reduced parameter size relative to its counterparts. Additionally, the FSG-Net-L achieved the highest F1 scores while possessing a median parameter size.
In the evolutionary history of CNNs, a noteworthy highlight is the dominance of U-Net [18] and its variants [14, 21, 20, 28, 13, 8, 22] as SOTA models in the field of medical image segmentation, especially for retinal vessel segmentation tasks on well-known datasets such as DRIVE, STARE, CHASE_DB1, and HRF. On the contrary, for clinical segmentation tasks on the CVC-clinic and Kvasir-SEG datasets, vision transformer models are employed [23, 2, 3, 9]. The primary distinction between retinal vessels and other clinical datasets lies in the level of feature intricacy required. Vision transform models exhibit some limitations in overcoming their constrained inductive bias. Meanwhile, our experiments underscore the ongoing relevance and effectiveness of attention mechanisms in addressing such challenges.
In this study, we aimed to propose a U-Net-based segmentation network that reflects the thin and elongated structural characteristics of retinal vessels, starting with the lower layers and gradually working our way up to the upper layers. Furthermore, based on the understanding that the guided filter [10] can function as a type of edge sharpening filter, the guided filter is located at the decoder after features from the encoder are merged so that full scale features can be used as input to the guided filters at each stage. Our network architecture improves both performance and computational efficiency simultaneously as shown in Figure 1. To ensure a fair comparison of the robustness of competing models, the training environment was maintained consistently.
The main contributions of this paper are summarized as follows: Firstly, enhancement of feature extraction through the proposal of a novel and efficient convolutional block. Secondly, proposing guided convolution blocks in the decoder, demonstrating performance enhancement as a result and enabling integration with various U-Net segmentation models. Thirdly, a comparison of various SOTA algorithms under a fixed experimental environment.
2 Method
2.1 Motivation
The guided filter was first introduced in [10] for image processing under the assumption that the guidance image and the filtered output have a locally linear relationship. The guided filter is formulated as:
$\hat{I_{i}}=a_{k}I_{i}+b_{k},\quad\forall i\in w_{k},$ (1)
where $a$, $b$ are linear coefficients, and $w$ is a local window. Here, we show why the guided filter can improve edge-like blood vessel segmentation performance. If we consider overlapping windows, the $i$-th pixel has several $a_{k}$ and $b_{k}$ values depending on the window size. Thus, the guided filter output of $I_{i}$ can be averaged as follows:
$\hat{I_{i}}=\frac{1}{|w|}\sum_{k\in{w_{i}}}\left(a_{k}I_{i}+b_{k}\right),\quad\forall i\in w_{k}.$ (2)
Now, (2) can be rewritten to have a more intuitive form. In the original paper of the guided filter [10], $b_{k}$ is computed as:
$b_{k}=\bar{p}_{k}-a_{k}\bar{I}_{k},$ (3)
where ()¯$\bar{\left(\cdot\right)}_{k}$ means average in $w_{k}$. Then, by putting Eq. (3) into Eq. (2), we have the following formulation:
$\hat{I}_{i}=\bar{a}_{i}(I_{i}-\bar{I}_{i})+\tilde{p}_{i},$ (4)
where ()~$\tilde{\left(\cdot\right)}_{k}$ denotes the average of the average, meaning $\tilde{p}_{i}=\frac{1}{|w|}\sum_{k\in w_{i}}{\bar{p}}_{k}$. The unsharp mask filtering is defined as $\hat{I}=\alpha\left(I-L(I)\right)+I$, where $L(\cdot)$ is a low-pass filter such as the Gaussian filter. Eq. (4) looks like unsharping mask filter, where sharpening mask $\left(I_{i}-\bar{I}_{i}\right)$ from the guidance image is added to the averaged target image. The filter strength is controlled by $\alpha$.
In their pioneering work [26], Zhang et al. adopted the guided filter in the segmentation network and suggested incorporating the attention map $M$ into the minimization problem to estimate $a$ and $b$:
$E(a_{k},b_{k})=\sum_{i\in{w_{k}}}\left(M_{i}^{2}\left(a_{k}{I_{d}}_{i}+b_{k}-g_{i}\right)^{2}+\mu a_{k}^{2}\right).$ (5)
In Eq. (5), the ${I_{d}}_{i}$ is the downsampled input feature from encoding parts, and the $g_{i}$ is the gating signal from upward path at the decoding parts. Eq. (5) implies that an improved attention map can lead to better solutions. In the following, we are going to propose a method that can generate an improved attention map with full-scale information.
2.2 Network architecture
Figure 2: Network architecture of the proposed FSG-Net.
As can be seen in Figure 2, the proposed network basically follows the U-Net architecture but has an additional guided convolution block after the feature representation network, unlike usual U-Net based structures. Contrary to the standard U-Net consisting of five stages, we have opted for four stages in FSG-Net based on the assumption that a wider receptive field is less critical for retinal vessel segmentation.
As described in the left part of FSG-Net, which is called the feature representation network, the input layer of the down-convolution is concatenated with a separate convolution from the other layers. These preserved features are then connected to the up-convolution layer. The deep bottleneck structure of the convolution block, indicated by the thick arrows, delivers a more enhanced feature representation. In the feature representation network, the key area to note is the feature merging section, indicated by the red dashed box. In this feature merging process, features from three paths—the current, upper, and lower stages—are concatenated and passed through a newly designed convolution block before being forwarded to the subsequent stages. This approach allows the integration and transmission of information across all scales.
In the right part of the model, namely the guided convolution block, compressed features reflecting full scale information are given as the current stage input to the guided residual module (GRM). Up-stage input from the higher stage is given as another input for guided filtering. After GRM, predictions are derived through convolution and activation. For training, a combination of the dice loss and BCE loss was employed to improve segmentation performance of the edge-like structures [7]:
$\begin{split}L_{\text{DS}}=\sum_{d=1}^{D}\sum_{i=0}^{S^{d}}\left(y_{i}\log\hat{y}_{i}^{d}+(1-y_{i})\log(1-\hat{y}_{i}^{d})\right),\end{split}$ (6)
$\begin{split}L_{Dice}=1-\sum_{d=1}^{D}\sum_{i=1}^{S^{d}}\frac{2\left(y_{i}\cdot\hat{y}_{i}^{(d)}\right)}{y_{i}+\hat{y}_{i}^{(d)}+\epsilon},\end{split}$ (7)
$\begin{split}L_{total}=L_{DS}+L_{Dice},\end{split}$ (8)
where $S$ denotes the indices and $L_{Dice}$ for Dice loss over pixels. For deep-supervision, $D$ represents the number of layers, the prediction at layer $d$, and weights for each layer, respectively.
Lastly, we can consider the scalability of the proposed architecture. The feature representation network can be replaced by other U-Net variants, enabling integration with the guided convolution block. For instance, we can consider integrating with FR-UNet [15], which proposes alternative methods for utilizing full-scale information.
2.3 Modernized convolution block
The standard U-Net has a fundamental structure comprising skip connections and double convolution blocks. Recent studies have shown that incorporating different CNN structures can significantly improve the performance of the original model [16, 9]. Inspired by the ConvNeXt [16], we have designed a convolutional block suitable for retinal vessel segmentation.
Figure 3: The evolutionary structure from (a): Depthwise residual block, (b): Inverted residual block to (c): The proposed convolution block.
Figure 4: Detailed structure of guided residual module (GRM)
Figure 3(a) and (b) represent the depthwise residual block and inverted residual block, respectively. Figure 3(c) shows the proposed convolution block. The structure of the block is an extension of the latest advancements in convolutional block development, characterized by features such as 1$\times$1 convolution, inverted bottleneck, and depth-wise convolution. Like other modernized convolution blocks, the proposed block employs a convolution with kernel size 2 and stride 2 in the first stage. By incorporating spatial and dimensional changes within the first block, we enable the design of a deep bottleneck stage in the down-convolution structure. This approach not only increases the information on the feature but also separates it from the bottleneck, allowing for more detailed feature representation. Furthermore, to maintain linearity, we utilize unique ReLU activation between each inverted residual block. For regularization purposes, we employ a learnable gamma parameter and apply a drop block before joining the identity block.
2.4 Guided convolution block
As shown in Figure 2, the guided convolution block consists of GRM to refine the input from the feature representation network and a convolution to output the prediction map. Figure 4 shows the overall process of GRM. The purpose of GRM is to refine the feature from the feature representation network at the same stage. In GRM, both the features at the current stage and at the higher stage are used to generate an attention map, and then the attention map is multiplied with the current feature to generate the output, which is described in the right part of Figure 4. Here we can expect an improved attention map $M$ can be estimated to solve Eq. (5) because the input feature from the feature representation network has full scale information. To further enhance the feature refinement, a residual block is introduced after the attention-guided filtering, as shown in the left part of Figure 4. By passing the features through the residual block, a stable map is generated for both the deep supervision and subsequent layers. Moreover, we incorporate a $1\times 1$ convolution to preserve the semantic information present in the feature maps.
3 Experiments
3.1 Training Techniques
In retinal vessel segmentation, the performance of studies is often determined by subtle gaps. We believe that these subtle gaps are highly influenced by the choice of hyperparameters and training/inference environment. To address this imbalance, we fixed all hyperparameters for training and inference. We empirically found that RandAugment [6] with a specific scale did not work well on medical datasets; therefore, we customized it to better suit our datasets. Training techniques include blur, color jitter, horizontal flip, perspective transformation, resize, crop, and CutMix [25].
Table 1: Train settings and hyper-parameters.
  Hyper-parameters Values
base lr 1e-3
lr scheduler Cosine annealing
lr scheduler warm-up epochs 20
lr scheduler cycle epochs 100
lr scheduler eta min 1e-5
early stop epochs 400
early stop metric F1 score
optimizer AdamW
optimizer momentum $\beta_{1},\beta_{2}$=0.9, 0.999
weight decay 0.05
criterion Dice + BCE
binary threshold 0.5
batch size 4
random crop 288
random blur prob=0.8
random jitter prob=0.8
random horizontal flip prob=0.5
random perspective prob=0.3
random random resize prob=0.8
CutMix n=1, prob=0.8
3.2 Implementation Details
Our experimental environment comprises an Intel Xeon Gold 5220 processor, a Tesla V100-SXM2-32GB GPU, Pytorch 1.13.1, and CUDA version 11.7. The inference time for FSG-Net-L was approximately $600ms$ for an input size of (608, 608) using a GPU-synchronized flow. To address as much variability as possible, we re-implemented comparison studies and integrated them into a single environment. To ensure experimental fairness, certain hyperparameters, including the framework, loss function, metric, data augmentation, and random seed, were fixed to measure the robustness of the model. Our training settings followed the hyperparameters in Table  1, used for segmentation tasks in ADE20K multiscale learning in ConvNeXt [16].
To evaluate the compared models under the same conditions, we prioritized the search for an optimized model in our training settings. For example, the learning rate can affect the gradient updating and training time, depending on the model’s parameter size and depth. Training for a predetermined number of epochs can result in diverging weights for heavy models and, conversely, for light models. Therefore, we chose the optimized model using an early stop based on cycles in the learning rate scheduler. To select the optimized model during the training step, we used the highest F1 score [27], with an early stop of 400 epochs. To balance exploitation and exploration in the learning parameters, we stack the batch to have more than two sets of mini-batches in one epoch with a learning rate scheduler. The detailed hyper-parameters are described in Table 1. With these experimental settings, the performance of the pure U-Net dramatically increased and even surpassed that of some recent studies, as shown in Table 2.
3.3 Datasets
The DRIVE dataset comprised 40 retinal images with a resolution of 565$\times$584 pixels, captured as part of a retinopathy screening study in the Netherlands. The STARE dataset comprises 20 retinal fundus images with a resolution of 700 × 605 pixels, and the CHASE_DB1 dataset includes 28 retinal images from schoolchildren with a resolution of 999$\times$960 pixels. Both the STARE and CHASE_DB1 datasets were manually annotated by two independent experts. We used the annotation of the first expert, named "Hoover A." in STARE and "1stHO" in CHASE_DB1, for our analysis. The HRF dataset comprises 45 images, equally divided into a 1:1:1 ratio of healthy patients, diabetic retinopaths, and glaucomatous patients, with a high resolution of 3504$\times$2336 pixels. To measure the performance of the models, it is necessary to divide the data into training and validation sets. As the retinal vessel segmentation dataset was relatively limited, we split the data into a 1:1 ratio of the training and validation sets. The DRIVE dataset was officially divided into training and validation sets, each containing 20 images. For the STARE, CHASE_DB1, and HRF datasets, we used the first half as training and the remaining half as valiation.
Table 2: Metric comparison of results on datasets.
  Architecture mIoU F1 score Acc AUC Sen MCC Rank Avg
DRIVE
U-Net 83.857 82.956 97.013 97.853 83.449 81.456 5
U-Net++ 81.228 79.564 96.524 96.271 77.802 77.830 13
U-Net3+ Deep 83.909 83.030 97.017 98.082 83.721 81.520 3.1
ResU-Net 83.862 82.953 97.021 97.766 83.226 81.453 5.8
ResU-Net++ 83.729 82.783 97.001 97.708 82.791 81.263 9.1
SAU-Net 83.368 82.334 96.925 97.616 82.311 80.782 11.2
DCASU-Net 83.743 82.808 96.996 97.838 83.080 81.290 7.8
AG-Net 83.176 82.111 96.882 97.628 82.155 80.540 11.8
AttU-Net 83.958 83.080 97.039 97.844 83.422 81.584 3.2
R2U-Net 83.555 82.580 96.952 97.879 82.961 81.038 8.8
ConvU-NeXt 83.800 82.882 97.012 97.835 83.019 81.367 7.2
FR-UNet 83.884 82.995 97.007 98.158 83.869 81.485 3.8
FSG-Net-L (ours) 84.068 83.229 97.042 98.235 84.207 81.731 1
STARE
U-Net 85.924 84.873 97.754 98.341 84.361 83.713 3.3
U-Net++ 81.514 79.061 97.022 94.479 75.539 77.764 12.8
U-Net3+ Deep 85.824 84.829 97.707 99.146 85.522 83.626 3.5
ResU-Net 85.964 84.872 97.767 98.050 83.997 83.726 3.5
ResU-Net++ 83.185 81.358 97.319 95.221 78.456 80.196 11.7
SAU-Net 85.158 84.061 97.604 97.784 84.015 82.885 7
DCASU-Net 84.423 83.064 97.454 98.362 83.583 81.771 8.8
AG-Net 84.811 83.766 97.565 98.403 83.347 82.516 7.7
AttU-Net 85.848 84.772 97.694 99.050 86.226 83.588 3.8
R2U-Net 83.727 81.786 97.468 93.457 77.048 80.810 11.3
ConvU-NeXt 85.339 84.186 97.658 97.866 83.401 82.998 6.7
FR-UNet 84.815 83.496 97.577 96.637 81.872 82.327 9.2
FSG-Net-L (ours) 86.118 85.100 97.746 98.967 86.608 83.958 1.7
CHASE_DB1
U-Net 82.065 80.159 97.404 99.368 85.370 79.002 6
U-Net++ 81.512 79.415 97.321 99.362 84.143 78.201 8.2
U-Net3+ Deep 82.489 80.697 97.483 99.506 85.740 79.558 2.7
ResU-Net 81.104 78.810 97.320 99.278 81.311 77.479 11.5
ResU-Net++ 73.276 67.411 95.966 96.223 67.782 65.593 13
SAU-Net 81.335 79.100 97.338 99.407 82.265 77.802 8.8
DCASU-Net 82.254 80.368 97.483 99.329 83.916 79.161 5.3
AG-Net 82.158 80.272 97.440 99.544 84.817 79.070 4.2
AttU-Net 82.562 80.742 97.546 99.430 83.907 79.537 3.3
R2U-Net 81.250 78.944 97.384 98.792 80.054 77.625 10.8
ConvU-NeXt 81.752 79.704 97.405 99.402 83.009 78.439 7
FR-UNet 81.330 79.170 97.269 99.544 84.744 77.944 7.7
FSG-Net-L (ours) 82.680 81.019 97.515 99.378 85.995 79.899 2.2
HRF
U-Net 82.291 81.329 97.093 98.571 82.756 79.868 4.8
U-Net++ 82.653 80.995 97.040 98.450 82.402 79.522 7.6
U-Net3+ Deep 83.006 81.445 97.124 98.427 82.436 79.997 4.1
ResU-Net 82.908 81.299 97.098 98.415 82.481 79.842 6.8
ResU-Net++ 77.008 73.096 96.032 95.312 70.587 71.251 12.8
SAU-Net 82.015 80.170 96.892 98.539 82.195 78.624 9.5
DCASU-Net 82.942 71.335 97.109 98.499 82.383 79.884 7.0
AG-Net 82.249 80.474 96.961 98.566 81.810 78.951 8.6
AttU-Net 83.017 81.448 97.113 98.558 82.830 79.996 2.6
R2U-Net 80.998 78.699 96.808 96.976 77.231 77.122 11.8
ConvU-NeXt 83.015 81.423 97.121 98.525 82.520 79.974 3.8
FR-UNet 82.431 80.709 97.011 98.294 81.677 79.218 9.5
FSG-Net-L (ours) 83.088 81.567 97.106 98.744 83.616 80.121 1.6
Figure 5: Qualitative evaluation on top-3 model structure on DRIVE validation set.
Figure 6: Inspection on output map at each stage $S_{i}$ on FSG-Net. The finer scale of map describes more detailed lines.
3.4 Results
In binary task evaluations, the Matthew correlation coefficient (MCC) is a powerful metric, as noted by Chicco et al. [4]. However, to avoid evaluations oriented towards a specific metric, we also report the average rank of each model, denoted as "Rank Avg" in Table 2. This average rank provides a measure of the stable performance of a model across different datasets. For example, FSG-Net, U-Net3+ [13], and AttU-Net [17] achieved high ranks in all three datasets, whereas ResU-Net and FR-UNet recorded inconsistent results across the three datasets. FSG-Net achieved the highest rank in terms of performance across all three datasets, recording dominant scores in mIoU, F1 score, and MCC, which is equivalent to a detailed expression of the segmentation map. Notably, FSG-Net outperformed the DRIVE dataset, recording SOTA in the F1 score and sensitivity. Figure 5 represents the predicted segmentation maps of the three best models in our evaluation metrics. The FSG-Net shows the best results, especially in segmenting thin vessels. FSG-Net incorporates deep supervision, which allows for the extraction of segmentation maps at each stage. Figure 6 shows the vessel detection results at various stages. It can be observed that as we move to lower levels, the performance in detecting thinner vessels improves. An interesting point is that even at coarse stage $S_{3}$, the detection of thicker vessels is generally well-maintained.
In the inference settings, we padded the original image with a multiple of 32 to preserve the flexible operation of specific models. Resizing the shape can lead to informational loss, which is critical in retinal vessel segmentation that requires high-fidelity maps. When measuring the metrics, we again removed the padding to generate a perfectly similar shape to the original image with no informational loss. With this unpadding trick, metrics that require true negatives and false negatives can be decreased compared with padded or resized images. Remarkably, the majority of models examined in our study exhibited superior performance compared to their original implementations in our training settings. For example, AG-Net achieved better performance on the DRIVE dataset (cIoU:69.71, Sen:82.16) in our environment than the results reported in the original paper (cIoU:69.65, Sen:81.00). Furthermore, U-Net, despite being an early model introduced over a decade ago, demonstrated robustness by achieving middle-range performance using only pure convolution layers.
3.5 Ablation study
To further understand the impact of the model capacity and structure on FSG-Net, we conducted ablation studies. In Table 3, we vary the depth of the down-convolution, base channel (Base_c) and structure. The F1 score of three datasets is used as a metric here. The results showed that even the smallest version of FSG-Net surpassed the other models. By comparing the scores in Table 2, the smallest FSG-Net with 1.17M parameters outperformed recent studies with an average rank of 5.2 across all metrics on the three datasets, compared to SAU-Net’s rank of 9.0 with a parameter size of 0.5M, DCASU-Net’s [24] rank of 7.0 with a parameter size of 2.6M, ConvU-NeXt’s rank of 6.7 with a parameter size of 3.5M, and FR-UNet’s rank of 6.6 with a parameter size of 7.4M.
Table 3: Ablation studies on model capacity and model structure
  model Base_c Depth Params (M) DRIVE STARE CHASE_DB1
FSG-Net-L 64 [3, 3, 9, 3] 18.32 83.229 85.100 81.019
FSG-Net-B 64 [2, 2, 6, 2] 14.46 83.191 84.934 79.658
FSG-Net-S 48 [3, 3, 9, 3] 10.33 83.145 84.917 80.529
FSG-Net-T 32 [3, 3, 9, 3] 4.61 83.098 84.698 79.982
FSG-Net-N 16 [3, 3, 9, 3] 1.17 82.904 84.692 79.431
       Structure
kernel size 3 to 7 35.77 83.006 84.200 80.044
AGF to AttU-Net like 19.61 83.115 84.671 79.333
down-conv to residual 45.60 82.893 82.734 75.765
w/o deep supervision 18.32 83.131 84.698 79.889
removing spatial attention 18.32 83.237 84.598 79.946
To optimize the network structure, we investigated the impact of increasing the kernel size, replacing the attention-guided filter with a simple attention block, and the effect of using pure residual blocks or removing the spatial attention. Our findings show that while increasing the kernel size expands the receptive field, it fails to improve performance in retinal vessel segmentation tasks that require detailed expression of the segmentation map. We found that the attention-guided filter proved to be robust and effective for this task. Additionally, we observed that using pure residual blocks in the deep bottleneck stage increases the parameter size and reduces computational efficiency without significantly improving feature extraction. Removing the spatial attention improved the performance in the DRIVE dataset; however, as the input size increased, the performance decreased proportionally.
Figure 7: Compared structure of feature merging in (a) FSG-Net, and (b) AG-Net.
4 Discussion
We previously emphasized the method of feature merging within the network architecture. This is because feature merging allows information from all stages to be mixed and provided as input to the guided filtering module. This approach can be compared to AG-Net, which also gathers features and provides them to the guided filter. Figure 7(a) illustrates the feature merging in the proposed method, while Figure 7(b) shows the corresponding part of AG-Net. Let us focus on the left path to the guided filtering modules in both cases. As can be seen in Figure 7(a), feature maps from three stages (current, upward, and downward) are concatenated and passed through the proposed convolution block. For AG-Net, only the features from the current and lower stages are merged for the attention-guided filter. In summary, our GRM gets feature information from all stages with a modernized convolution, while the AG-Net gets it from current and lower stages with conventional convolutions. For this reason, we were able to achieve superior performance with one fewer stage than AG-Net.
The next discussion point concerns the experimental datasets. DRIVE, STARE, and Chase_DB1 are representative datasets for retinal vessel segmentation. Models that improve the architecture of layers and attention mechanisms continue to emerge, but the number of images in such datasets is limited. Although this paper focuses on retinal vessel segmentation, it is necessary to apply the proposed model to large databases from other fields to verify its variability and scalability. In our future work, we can consider expanding our application areas to include crack detection, facial wrinkle detection, and general medical imaging.
5 Conclusion
In this study, we presented a full-scale representation guided network for retinal vessel segmentation that achieves SOTA on the DRIVE dataset. A modern convolutional block tailored for retinal vessel segmentation was designed, and the guided convolution block maximized performance using full-scale information from the feature representation network. The proposed guided convolution block is compatible with any U-Net architecture, offering scalability for similar tasks across various domains. We encourage future research to use the FSG-Net architecture and our experimental settings to ensure reproducibility and scalability in similar tasks.