### v2

# Full-scale Representation Guided Network for Retinal Vessel Segmentation

Abstract
The U-Net architecture and its variants have remained state-of-the-art (SOTA) for retinal vessel segmentation over the past decade. In this study, we introduce a Full-Scale Guided Network (FSG-Net), where a novel feature representation module using modernized convolution blocks effectively captures full-scale structural information, while a guided convolution block subsequently refines this information. Specifically, we introduce an attention-guided filter within the guided convolution block, leveraging its similarity to unsharp masking to enhance fine vascular structures. Passing full-scale information to the attention block facilitates the generation of more contextually relevant attention maps, which are then passed to the attention-guided filter, providing further refinement to the segmentation performance. The structure preceding the guided convolution block can be replaced by any U-Net variant, ensuring flexibility and scalability across various segmentation tasks. For a fair comparison, we re-implemented recent studies available in public repositories to evaluate their scalability and reproducibility. Our experiments demonstrate that, despite its compact architecture, FSG-Net delivers performance competitive with SOTA methods across multiple public datasets. Ablation studies further demonstrate that each proposed component meaningfully contributes to this competitive performance. Our code is available on https://github.com/ZombaSY/FSG-Net-pytorch.
Introduction
Convolutional neural networks (CNNs) have seen significant improvements in performance and optimization since the 2010s. The introduction of hardware acceleration using GPUs, ReLU activation function, and residual block [He_2016_CVPR, https://doi.org/10.48550/arxiv.1802.06955] has enabled smooth back-propagation in deep neural network architectures. Research in this field has focused on finding a balance between computational efficiency, parameter size, code scalability, and fidelity. Depthwise separable convolution [Chollet_2017_CVPR] and squeeze-and-excitation [Hu_2018_CVPR] have been particularly influential in this regard. The inverted residual block [Sandler_2018_CVPR] achieved higher fidelity with optimized computational efficiency and a smaller parameter size compared to ResNet.
Figure 1: F1 scores of compared networks on the DRIVE dataset, measured against the validation dataset comprising zero-padded images of resolution $\times608$. Among the considered architectures, FSG-Net-T achieved a superior F1 score compared to competitive models while maintaining a reduced parameter size relative to its counterparts. Additionally, the FSG-Net achieved the highest F1 scores while possessing a median parameter size.
In the evolutionary history of CNNs, a noteworthy highlight is the dominance of U-Net [10.1007/978-3-319-24574-4_28] and its variants as SOTA models in the field of medical image segmentation across a wide spectrum of segmentation tasks [8959021, Tang_2022_CVPR, 9446143, 10.1007/978-3-030-00889-5_1, 9053405, 9413346, TIAN2020117, jin2022semi, jin2024inter, jin2021cascade, jin2021domain, tan2024deep, ding2024rcar]. In particular, retinal-vessel segmentation benefits from well-curated public datasets such as DRIVE, STARE, CHASE_DB1, and HRF, on which the latest CNN-based methods are already systematically benchmarked and firmly established as strong baselines. In [qin2024review], traditional and the recent deep learning based algorithms for retinal vessel segmentation are comprehensively reviewed.
On the contrary, for clinical segmentation tasks on the CVC-clinic and Kvasir-SEG datasets, vision transformer (ViT) models are employed [xie2021segformer, 10.1007/978-3-031-25066-8_9, chen2021transunet, HAN2022109512]. The primary distinction between retinal vessels and other clinical datasets lies in the level of feature intricacy required. ViT models exhibit some limitations in overcoming their constrained inductive bias. Meanwhile, our experiments underscore the ongoing relevance and effectiveness of attention mechanisms in addressing such challenges.
In this study, we aimed to propose a U-Net-based segmentation network that explicitly captures and preserves the thin and elongated structural characteristics of retinal vessels by leveraging a novel, full-scale feature representation strategy from the early layers upward. Furthermore, we introduced an attention-guided filtering approach inspired by guided filters [6319316], strategically embedded in the decoder following encoder-decoder feature merging. This design allows the guided filter to perform adaptive edge enhancement, specifically utilizing comprehensive, full-scale structural information at each decoding stage. Our network architecture improves both performance and computational efficiency simultaneously, as shown in Fig. 1. To ensure a fair comparison of the robustness of competing models, the training environment was maintained consistently.
The main contributions of this paper are summarized as follows: Firstly, a novel convolutional block specifically designed for retinal vessel segmentation, capable of extracting comprehensive, full-scale structural features with computational efficiency. Secondly, an attention-guided convolutional block integrated into the decoder, enhancing segmentation accuracy through adaptive refinement, and offering flexibility to be combined seamlessly with various U-Net architectures. Thirdly, a thorough comparison and reproducibility study of multiple SOTA algorithms, conducted under strictly controlled experimental conditions.
Related work
Guided filter
The guided filter was first introduced in [6319316] for image processing under the assumption that the guidance image and the filtered output have a locally linear relationship. The guided filter is formulated as:
$\hat{I_{i}}=a_{k}I_{i}+b_{k},\quad\forall i\in\mathrm{w}_{k},$ (1)
where $a_{k}$, $b_{k}$ are linear coefficients and assumed to be constants in the local window $\mathrm{w}_{k}$. $\hat{I}_{i}$ is the output of the guided filter with input $I_{i}$. Here, we show why the guided filter can improve edge-like blood vessel segmentation performance. If we consider overlapping windows, the $i$-th pixel has several $a_{k}$ and $b_{k}$ values depending on the window size. Thus, the guided filter output of $I_{i}$ can be averaged as follows:
$\hat{I_{i}}=\frac{1}{|w|}\sum_{k\in{w_{i}}}\left(a_{k}I_{i}+b_{k}\right),\quad\forall i\in\mathrm{w}_{k}.$ (2)
The guided filter smooths images while preserving edge information, making it widely used in various image processing tasks such as image detail enhancement, image denoising, image preprocessing, and image dehazing.
Multiscale architectures preserving spatial detail
In this study, we focus on two representative types of multiscale architectures for image segmentation: full-resolution preserving and full-scale aggregation approaches. In the full-resolution preserving category, HRNet maintains high-resolution representations by processing multi-resolution branches in parallel throughout the network [Sun_2019_CVPR]. Similarly, FR-UNet adopts a parallel full-resolution pathway while preserving the U-Net structure [9815506]. In the full-scale aggregation category, UNet3+ extends the skip-connection scheme of the original U-Net by enabling each decoder stage to simultaneously receive and fuse feature maps from all encoder and decoder stages [9053405].
Deep supervision
Deep supervision has recently become a prevalent strategy in various segmentation tasks, particularly within medical image analysis, to enhance model performance [10.1007/978-3-030-00889-5_1, wang2019deeply, reiss2021every]. In general, deep supervision refers to computing multiple loss functions at different intermediate layers of the neural network, enabling more stable gradient propagation throughout the training process. This approach has proven beneficial in reducing issues such as vanishing gradients, thus improving segmentation accuracy.
Several well-known variations of U-Net architectures, including M-Net [fu2018joint], AG-Net [10.1007/978-3-030-32239-7_88], and U-Net3+ [9053405], have successfully applied deep supervision and demonstrated better segmentation results compared to architectures utilizing a single loss function. Specifically, for retinal vessel segmentation, M-Net and AG-Net calculated multiple losses from each of the decoders, except the deepest one. On the other hand, U-Net3+ utilized losses from all decoder stages and employed them to train a segmentation model targeting liver and spleen structures in CT images.
Method
Motivation
Now, (2) can be rewritten to have a more intuitive form. In the original paper of the guided filter [6319316], $b_{k}$ is computed as:
$b_{k}=\bar{p}_{k}-a_{k}\mu_{k},$ (3)
where $\bar{\left(\cdot\right)}_{k}$ means average in a window $\mathrm{w}_{k}$ and $\mu_{k}$ is the mean of $I_{k}$. Then, by putting Eq. (3) into Eq. (2), we have the following formulation:
$\hat{I}_{i}=\bar{a}_{i}(I_{i}-\mu_{i})+\tilde{p}_{i},$ (4)
where $\tilde{\left(\cdot\right)}_{k}$ denotes the average of the average, meaning $\tilde{p}_{i}=\frac{1}{|w|}\sum_{k\in w_{i}}{\bar{p}}_{k}$. The unsharp masking is defined as follows:
$\hat{I}_{i}=\alpha\left(I_{i}-\sum_{k\in\mathrm{w}_{i}}w_{ik}I_{k}\right)+I_{i},$ (5)
where $w_{ik}$ denotes the normalized Gaussian weight centered at pixel $i$, evaluated at neighboring pixel $k$. Eq. (4) looks like Eq. (5), where sharpening mask $\left(I_{i}-\bar{I}_{i}\right)$ from the guidance image is added to the averaged target image. The filter strength is controlled by $\alpha$.
In their pioneering work [10.1007/978-3-030-32239-7_88], Zhang et al. adopted the guided filter in the segmentation network and suggested incorporating the attention map $M$ into the energy minimization problem to estimate $a$ and $b$:
$\mathcal{E}(a_{k},b_{k})=\sum_{i\in\mathrm{w}_{k}}\left(M_{i}^{2}\left(a_{k}{I_{d}}_{i}+b_{k}-g_{i}\right)^{2}+\epsilon a_{k}^{2}\right).$ (6)
In Eq. (6), the ${I_{d}}_{i}$ is the downsampled input feature from encoding parts, where $g_{i}$ denotes the gating signal from upward path at the decoding parts. Eq. (6) implies that an improved attention map can lead to better solutions. In the following, we are going to propose a method that can generate an improved attention map with full-scale information.
Network architecture
Figure 2: Network architecture of the proposed FSG-Net. In the feature representation network (left), down-convolution layers concatenate separately extracted features, connected via deep bottleneck structures to up-convolution layers. The red dashed boxes highlight feature merging from multiple scales. The guided convolution block (right) employs Guided Residual Modules (GRM) to refine multi-scale features. $S_{i}$ indicates each stage, and $y_{i}$ represents labels, where $y_{2}$ and $y_{3}$ are downsampled from the original label $y_{1}$ by interpolation.
As can be seen in Fig. 2, the proposed network basically follows the U-Net architecture but has an additional guided convolution block after the feature representation network, unlike usual U-Net based structures. Contrary to the standard U-Net consisting of five stages, we have opted for four stages in FSG-Net based on the assumption that a wider receptive field is less critical for retinal vessel segmentation.
As illustrated on the left side of the FSG-Net architecture, the input to each down-convolution block, which is indicated by the bold green arrows in the figure, is augmented by concatenating features derived from an auxiliary convolutional path. These preserved features are subsequently forwarded to the corresponding up-convolution layers via skip connections, facilitating richer multi-scale feature reuse. The convolutional blocks within this stage adopt a deep bottleneck structure, which enhances the network’s representational capacity while maintaining computational efficiency. Furthermore, a lightweight spatial attention, inspired by the Convolutional Block Attention Module (CBAM) [Woo_2018_ECCV], is integrated into the bottleneck stage to refine spatial feature encoding with minimal parameter overhead.
In the feature representation network, the key area to note is the feature merging section, indicated by the red dashed box. In this feature merging process, features from three paths—the current, upper, and lower stages—are concatenated and passed through a newly designed convolution block before being forwarded to the subsequent stages. This approach allows the integration and transmission of information across all scales. In this study, we refer to such integrated multi-scale information as full-scale information. The advantages of gathering the entire information in this manner will be further discussed in the Results section.
In the right part of the model, namely the guided convolution block, compressed features reflecting full scale information are given as the current stage input to the guided residual module (GRM). Up-stage input from the higher stage is given as another input for guided filtering. After GRM, predictions are derived through convolution and activation. For training, a combination of the BCE loss and Dice loss with deep supervision was employed to improve segmentation performance of the edge-like structures [rs15061530]:
$L_{\text{DS}}=\sum_{d=1}^{D}\alpha_{d}\cdot\left(L_{\text{BCE}}^{(d)}+\lambda\cdot L_{\text{Dice}}^{(d)}\right),$ (7)
$\text{where}\quad L_{\text{BCE}}^{(d)}=-\frac{1}{S_{d}}\sum_{i=1}^{S_{d}}\left(y_{i,d}\log\hat{y}_{i,d}+(1-y_{i,d})\log(1-\hat{y}_{i,d}\right),$ (8)
$\text{and}\quad L_{\text{Dice}}^{(d)}=1-\frac{2\sum_{i=1}^{S_{d}}y_{i,d}\hat{y}_{i,d}+\epsilon}{\sum_{i=1}^{S_{d}}y_{i,d}+\sum_{i=1}^{S_{d}}\hat{y}_{i,d}+\epsilon}.$ (9)
In Eq. (7), $D$, $\alpha_{d}$ represent the number of prediction layers and weights for each prediction layer $d$, respectively. In Eq. (8) and Eq. (9), $S_{d}$ denotes the number of samples at the $d$-th prediction layer and $y_{i,d}$ and $\hat{y}_{i,d}$ represent target and prediction values at $i$-th index and at $d$-th prediction level, respectively.
Lastly, we can consider the scalability of the proposed architecture. The feature representation network can be replaced by other U-Net variants, enabling integration with the guided convolution block. For instance, we can consider integrating with FR-UNet [9815506], which proposes alternative methods for utilizing full-scale information.
Modernized convolution block
The standard U-Net has a fundamental structure comprising skip connections and double convolution blocks. Recent studies have shown that incorporating different CNN structures can significantly improve the performance of the original model [liu2022convnet, HAN2022109512]. Inspired by the ConvNeXt [liu2022convnet], we have designed a convolutional block suitable for retinal vessel segmentation.
Figure 3: The evolutionary structure from (a): Depthwise residual block, (b): Inverted residual block to (c): The proposed convolution block.
Fig. 3(a) and (b) represent the depthwise residual block and inverted residual block, respectively. Fig. 3(c) shows the proposed convolution block. The structure of the block is an extension of the latest advancements in convolutional block development, characterized by features such as $\times1$ convolution, inverted bottleneck, and depth-wise convolution. Like other modernized convolution blocks, the proposed block employs a convolution with kernel size 2 and stride 2 in the first stage. By incorporating spatial and dimensional changes within the first block, we enable the design of a deep bottleneck stage in the down-convolution structure. This approach not only increases the information on the feature but also separates it from the bottleneck, allowing for more detailed feature representation. Furthermore, to maintain linearity, we utilize unique ReLU activation between each inverted residual block. For regularization purposes, we employ a learnable gamma parameter and apply a drop block before joining the identity block.
Guided convolution block
As shown in Fig. 2, the guided convolution block consists of GRM to refine the input from the feature representation network and a convolution to output the prediction map. Fig. 4 shows the overall process of GRM.
The purpose of GRM is to refine the feature from the feature representation network at the same stage. In GRM, both the features at the current stage and at the higher stage are used to generate an attention map, and then the attention map is multiplied with the current feature to generate the output, which is described in the right part of Fig. 4. Here we can expect an improved attention map $M$ can be estimated to solve Eq. (6) because the input feature from the feature representation network has full scale information. To further enhance the feature refinement, a residual block is introduced after the attention-guided filtering, as shown in the left part of Fig. 4. By passing the features through the residual block, a stable map is generated for both the deep supervision and subsequent layers. Moreover, we incorporate a $1\times 1$ convolution to preserve the semantic information present in the feature maps.
Figure 4: Detailed structure of guided residual module (GRM)
Experiments
Training Techniques
In retinal vessel segmentation, the performance of studies is often determined by subtle gaps. We believe that these subtle gaps are highly influenced by the choice of hyperparameters and training/inference environment. To address this imbalance, we fixed all hyperparameters for training and inference. We empirically found that RandAugment [Cubuk_2020_CVPR_Workshops] with a specific scale did not work well on medical datasets; therefore, we customized it to better suit our datasets. Training techniques include blur, color jitter, horizontal flip, perspective transformation, resize, crop, and CutMix [Yun_2019_ICCV].
Implementation Details
[…]
Table 1: Train settings and hyper-parameters.
  Hyper-parameters Values
base lr 1e-3
lr scheduler Linear warm-up, Cosine annealing
lr scheduler warm-up epochs 20
lr scheduler cycle epochs 100
lr scheduler eta min 1e-5
early stop epochs 400
early stop metric F1 score
optimizer AdamW
optimizer momentum $\beta_{1},\beta_{2}$=0.9, 0.999
weight decay 0.05
criterion Dice + BCE
binary threshold 0.5
batch size 4
center padded shape (dataset) (D=608, S=704, C=1024, H=1344)
random crop 288
[…]
random jitter b=0.2, c=0.2, s=0.2, h=0.1, prob=0.8
random horizontal flip prob=0.5
random perspective s=0.3, prob=0.3
random random resize s=[0.5, 2.0], prob=0.8
CutMix n=1, prob=0.8
To evaluate the compared models under the same conditions, we prioritized the search for an optimized model in our training settings. For example, the learning rate can affect the gradient updating and training time, depending on the model’s parameter size and depth. Training for a predetermined number of epochs can result in diverging weights for heavy models and, conversely, for light models. Therefore, we chose the optimized model using an early stop based on cycles in the learning rate scheduler. To select the optimized model during the training step, we used the highest F1 score [8309343], with an early stop of 400 epochs. To balance exploitation and exploration in the learning parameters, we stack the batch to have more than two sets of mini-batches in one epoch with a learning rate scheduler. The detailed hyper-parameters are described in Table 1. With these experimental settings, the performance of the pure U-Net dramatically increased and even surpassed that of some recent studies, as shown in Table 2.
Datasets
[…]
Table 2: Comparison of segmentation performance for CNN-based networks.
Architecture
mIoU
F1 score
Acc
AUC
Sen
MCC
Rank Avg
DRIVE
U-Net 83.857 82.956 97.013 97.853 83.449 81.456 5.7
U-Net++ 81.228 79.564 96.524 96.271 77.802 77.830 14.0
U-Net3+ Deep 83.909 83.030 97.017 98.082 83.721 81.520 3.5
ResU-Net 83.862 82.953 97.021 97.766 83.226 81.453 6.5
ResU-Net++ 83.729 82.783 97.001 97.708 82.791 81.263 10.2
SAU-Net 83.368 82.334 96.925 97.616 82.311 80.782 12.2
DCASU-Net 83.743 82.808 96.996 97.838 83.080 81.290 8.7
AG-Net 83.176 82.111 96.882 97.628 82.155 80.540 12.8
AttU-Net 83.958 83.080 97.039 97.844 83.422 81.584 3.5
R2U-Net 83.555 82.580 96.952 97.879 82.961 81.038 9.7
ConvU-NeXt 83.800 82.882 97.012 97.835 83.019 81.367 7.8
FR-UNet 83.884 82.995 97.007 98.158 83.869 81.485 4.3
HRNet 83.938 83.829 97.325 97.860 82.963 81.506 5.0
FSG-Net (ours) 84.068 83.229 97.042 98.235 84.207 81.731 1.2
STARE
U-Net 85.924 84.873 97.754 98.341 84.361 83.713 3.7
U-Net++ 81.514 79.061 97.022 94.479 75.539 77.764 13.8
U-Net3+ Deep 85.824 84.829 97.707 99.146 85.522 83.626 3.7
ResU-Net 85.964 84.872 97.767 98.050 83.997 83.726 4.0
ResU-Net++ 83.185 81.358 97.319 95.221 78.456 80.196 12.7
SAU-Net 85.158 84.061 97.604 97.784 84.015 82.885 8.0
DCASU-Net 84.423 83.064 97.454 98.362 83.583 81.771 9.7
AG-Net 84.811 83.766 97.565 98.403 83.347 82.516 8.5
AttU-Net 85.848 84.772 97.694 99.050 86.226 83.588 4.0
R2U-Net 83.727 81.786 97.468 93.457 77.048 80.810 12.3
ConvU-NeXt 85.339 84.186 97.658 97.866 83.401 82.998 7.5
FR-UNet 84.815 83.496 97.577 96.637 81.872 82.327 10.2
HRNet 85.555 84.165 97.982 98.190 84.604 83.197 5.2
FSG-Net (ours) 86.118 85.100 97.746 98.967 86.608 83.958 1.8
CHASE_DB1
U-Net 82.065 80.159 97.404 99.368 85.370 79.002 7.0
U-Net++ 81.512 79.415 97.321 99.362 84.143 78.201 9.2
U-Net3+ Deep 82.489 80.697 97.483 99.506 85.740 79.558 3.7
ResU-Net 81.104 78.810 97.320 99.278 81.311 77.479 12.5
ResU-Net++ 73.276 67.411 95.966 96.223 67.782 65.593 14.0
SAU-Net 81.335 79.100 97.338 99.407 82.265 77.802 9.8
DCASU-Net 82.254 80.368 97.483 99.329 83.916 79.161 6.3
AG-Net 82.158 80.272 97.440 99.544 84.817 79.070 5.2
AttU-Net 82.562 80.742 97.546 99.430 83.907 79.537 4.3
R2U-Net 81.250 78.944 97.384 98.792 80.054 77.625 11.8
ConvU-NeXt 81.752 79.704 97.405 99.402 83.009 78.439 8
FR-UNet 81.330 79.170 97.269 99.544 84.744 77.944 8.7
HRNet 82.849 81.021 97.744 99.650 85.750 79.987 1.2
FSG-Net (ours) 82.680 81.019 97.515 99.378 85.995 79.899 3.0
HRF
U-Net 82.291 81.329 97.093 98.571 82.756 79.868 4.8
U-Net++ 82.653 80.995 97.040 98.450 82.402 79.522 7.8
U-Net3+ Deep 83.006 81.445 97.124 98.427 82.436 79.997 4.4
ResU-Net 82.908 81.299 97.098 98.415 82.481 79.842 7.0
ResU-Net++ 77.008 73.096 96.032 95.312 70.587 71.251 13.8
SAU-Net 82.015 80.170 96.892 98.539 82.195 78.624 10.3
DCASU-Net 82.942 71.335 97.109 98.499 82.383 79.884 7.3
AG-Net 82.249 80.474 96.961 98.566 81.810 78.951 8.7
AttU-Net 83.017 81.448 97.113 98.558 82.830 79.996 2.8
R2U-Net 80.998 78.699 96.808 96.976 77.231 77.122 12.8
ConvU-NeXt 83.015 81.423 97.121 98.525 82.520 79.974 4.0
FR-UNet 82.431 80.709 97.011 98.294 81.677 79.218 9.7
HRNet 82.067 80.226 96.944 98.563 80.957 78.668 9.8
FSG-Net (ours) 83.088 81.567 97.106 98.744 83.616 80.121 1.6
Results
In binary task evaluations, the Matthew correlation coefficient (MCC) is a powerful metric, as noted by Chicco et al. [Chicco2021TheBO]. However, to avoid evaluations oriented towards a specific metric, we also report the average rank of each model, denoted as "Rank Avg" in Table 2. This average rank provides a measure of the stable performance of a model across different datasets. For example, FSG-Net, U-Net3+ [9053405], and AttU-Net [DBLP:journals/corr/abs-1804-03999] achieved high ranks in all four datasets, whereas ResU-Net, FR-UNet and HRNet [Sun_2019_CVPR] recorded inconsistent results across the three datasets. FSG-Net consistently demonstrated top-tier performance across all four datasets, recording dominant scores in mIoU, F1 score, and MCC, which is equivalent to a detailed expression of the segmentation map. Notably, FSG-Net outperformed previous methods on the DRIVE dataset, achieving SOTA performance in F1 score and sensitivity.
We conducted comparative experiments with existing architectures that aim to preserve full-resolution features, namely HRNet and FR-UNet. As shown in Table 2, HRNet achieves the best performance on the CHASE_DB1 dataset and also records the highest accuracy on the DRIVE and STARE datasets. However, its performance noticeably drops on the HRF dataset, and notably, HRNet contains 65.81M parameters, which significantly exceeds the model capacities of all other compared architectures that maintain parameter counts below 30M. FR-UNet performs well on the DRIVE dataset but shows relatively moderate results on the remaining datasets. These comparisons with high-resolution preserving models demonstrate the performance stability and robustness of FSG-Net across diverse datasets, validating the effectiveness of utilizing full-scale information at multiple decoding stages.
Table 3: Performance comparison of ViT-based models and FSG-Net on four retinal vessel segmentation datasets.
Architecture
mIoU
F1 score
Acc
AUC
Sen
MCC
DRIVE
Swin 78.758 76.040 96.296 97.203 75.681 74.171
MaxViT-512 79.702 77.738 96.050 98.452 79.097 75.693
FSG-Net (ours) 18.32 84.068 83.229 97.042 98.235 84.207 81.731
STARE
Swin 79.354 75.900 96.958 97.759 75.186 74.356
MaxViT-512 81.072 78.898 96.816 97.865 79.667 77.357
FSG-Net (ours) 86.118 85.100 97.746 98.967 86.608 83.958
CHASE_DB1
Swin 77.708 73.850 96.966 99.326 76.569 72.347
MaxViT-512 79.322 76.405 96.978 99.494 79.878 74.909
FSG-Net (ours) 82.680 81.019 97.515 99.378 85.995 79.899
HRF
Swin 78.739 74.771 97.420 98.022 74.831 73.548
MaxViT-512 74.980 70.617 95.346 98.645 72.600 68.204
FSG-Net (ours) 83.088 81.567 97.106 98.744 83.616 80.121
We have also conducted experiments to assess the suitability of ViT-based models for the task of retinal vessel segmentation. As noted in the introduction, the thin and elongated structure of retinal vessels poses specific challenges, particularly for models that rely heavily on global representations such as pure ViT. To explore this further, we evaluated Swin-T [Liu_2021_ICCV], which introduces a hierarchical representation with shifted window attention, and MaxViT-T-512 [10.1007/978-3-031-20053-3_27], a hybrid model that combines convolution and attention mechanisms and was evaluated using the UPerNet [Xiao_2018_ECCV] decoding head. As shown in Table 3, the ViT-based models in our setting yielded lower overall performance compared to FSG-Net, particularly in metrics such as F1 score and sensitivity. Swin-T and MaxViT-T-512 contain 58.91M and 59.60M parameters, respectively, yet FSG-Net achieves substantially better performance despite operating with a more parameter-efficient design.
Nonetheless, recent advances in DETR-like architectures [10.1007/978-3-030-58452-8_13] have introduced object query mechanisms that demonstrate strong potential for enhancing small object detection. These efforts reflect ongoing attempts to overcome the limitations traditionally associated with representing fine, low-saliency structures using transformer-based models. Although not included in our experiments, TCDDU-Net [lv2024tcddu], which combines a dual-path U-Net architecture with a Swin backbone, has been reported to outperform several CNN-based methods in terms of quantitative metrics and to effectively segment peripheral vessels. Although not directly comparable due to different experimental settings, it is notable that TCDDU-Net achieved an F1 score of 82.65 on the DRIVE dataset, which is slightly lower than the 84.068 recorded by our proposed FSG-Net. These research findings suggest that future research may benefit from further exploring hybrid architectures or improving transformer-based designs to better handle fine-grained and low-saliency structures such as retinal vessels.
Fig. 5 represents the predicted segmentation maps of the three best models in our evaluation metrics. The FSG-Net shows the best results, especially in segmenting thin vessels. Fig. 6 shows prediction results of FSG-Net obtained through deep supervision from the decoder stages indicated in Fig. 2. As shown in Fig. 6, intermediate predictions may not directly provide fine-grained details, but they can improve the final segmentation results due to their accurate semantic-level information and the capability to capture global context. The quantitative performance benefits of such deep supervision is further discussed in the ablation study.
Figure 5: Qualitative comparison of the top-3 performing models on the DRIVE validation set.
Figure 6: Visualization of predictions at each stage $S_{i}$ of FSG-Net. Here, $S_{i}$ represents each decoder stage, and a larger value of $i$ indicates a deeper decoder as shown in Fig. 2. The first four rows show the entire prediction results from various datasets, while the last three rows present magnified views of these predictions.
In the inference settings, we padded the original image with a multiple of 32 to preserve the flexible operation of specific models. Resizing the shape can lead to informational loss, which is critical in retinal vessel segmentation that requires high-fidelity maps. When measuring the metrics, we again removed the padding to generate a perfectly similar shape to the original image with no informational loss. With this unpadding trick, metrics that require true negatives and false negatives can be decreased compared with padded or resized images. Remarkably, the majority of models examined in our study exhibited superior performance compared to their original implementations in our training settings. For example, AG-Net achieved better performance on the DRIVE dataset (cIoU:69.71, Sen:82.16) in our environment than the results reported in the original paper (cIoU:69.65, Sen:81.00). Furthermore, U-Net, despite being an early model introduced over a decade ago, demonstrated robustness by achieving middle-range performance using only pure convolution layers.
Table 4: Cross-domain evaluation across datasets, DRIVE (D), CHASE_DB1 (C), STARE (S), HRF (H). $D\rightarrow C$ indicates that the model was trained on the (D) dataset and evaluated on the (C) dataset.
Dataset
mIoU
F1 score
Acc
AUC
Sen
MCC
$D\rightarrow C$ 79.76 (-2.92) 77.31 (-3.70) 96.46 (-1.05) 99.01 (-0.36) 87.84 (+1.85) 76.14 (-3.75)
$D\rightarrow S$ 82.14 (-3.97) 80.29 (-4.81) 96.95 (-0.79) 97.77 (-1.19) 82.93 (-3.67) 78.90 (-5.05)
$D\rightarrow H$ 79.24 (-3.84) 76.45 (-5.11) 96.21 (-0.89) 98.92 (+0.18) 81.93 (-1.68) 74.75 (-5.37)
$C\rightarrow D$ 78.38 (-5.68) 75.56 (-7.66) 96.25 (-0.79) 88.56 (-9.67) 66.98 (-17.2) 74.44 (-7.29)
$C\rightarrow S$ 81.51 (-4.60) 79.54 (-5.56) 96.98 (-0.76) 96.34 (-2.62) 77.52 (-9.08) 78.09 (-5.86)
$C\rightarrow H$ 80.25 (-2.83) 77.77 (-3.79) 96.57 (-0.53) 97.96 (-0.78) 79.86 (-3.75) 76.13 (-3.99)
$S\rightarrow D$ 80.70 (-3.36) 78.82 (-4.40) 96.52 (-0.52) 94.21 (-4.02) 74.84 (-9.36) 77.22 (-4.51)
$S\rightarrow C$ 79.87 (-2.81) 77.50 (-3.51) 96.56 (-0.95) 99.20 (-0.17) 86.04 (+0.05) 76.16 (-3.73)
$S\rightarrow H$ 78.38 (-4.70) 76.01 (-5.55) 96.08 (-1.02) 98.79 (+0.05) 82.49 (-1.12) 74.33 (-5.79)
$H\rightarrow D$ 79.13 (-4.93) 76.64 (-6.58) 96.27 (-0.77) 92.27 (-5.96) 70.79 (-13.4) 75.07 (-6.66)
$H\rightarrow C$ 81.80 (-0.88) 79.88 (-1.13) 97.10 (-0.41) 99.59 (+0.22) 83.69 (-2.30) 78.46 (-1.43)
$H\rightarrow S$ 82.18 (-3.93) 80.15 (-4.95) 97.15 (-0.59) 95.06 (-3.90) 76.81 (-9.79) 78.89 (-5.06)
To investigate the domain generalization capability of our model in retinal vessel segmentation, we conducted cross-domain validation experiments as presented in Table 4. Specifically, we trained the model on the DRIVE dataset and evaluated its performance on different unseen datasets, namely CHASE_DB1 (D $\rightarrow$ C), STARE (D $\rightarrow$ S), and HRF (D $\rightarrow$ H). As expected, performance decreased across most evaluation metrics when the model was tested on datasets different from the training domain. For example, in the D→C scenario, we observed reductions in mIoU (-2.92), F1 score (-3.70), and MCC (-3.75), although sensitivity showed a slight improvement (+1.85). Similar observations were made in other domain pairs, notably D→S and D→H, reflecting the inherent challenges posed by domain shifts such as differences in imaging modalities, resolutions, and contrast distributions.
However, despite this performance drop, the model still maintained reasonable predictive capability, indicating its robustness and capacity to learn transferable, domain-invariant features. These findings underscore the importance of enhancing domain generalization by exploring additional techniques, such as domain augmentation, pre-training strategies, linear probing, or knowledge distillation, to further improve the model’s adaptability to unseen domains.
Ablation study
Table 5: Ablation study on model capacity.
model
Base_c
Depth
Params(M)
DRIVE
STARE
CHASE_DB1
FSG-Net-L (proposed) 64 [3, 3, 9, 3] 18.32 83.229 85.100 81.019
FSG-Net-B 64 [2, 2, 6, 2] 14.46 83.191 84.934 79.658
FSG-Net-S 48 [3, 3, 9, 3] 10.33 83.145 84.917 80.529
FSG-Net-T 32 [3, 3, 9, 3] 4.61 83.098 84.698 79.982
FSG-Net-N 16 [3, 3, 9, 3] 1.17 82.904 84.692 79.431
To further understand the impact of the model capacity and structure on FSG-Net, we conducted ablation studies. In Table 5, we vary the depth of the down-convolution, base channel (Base_c) and structure. The F1 score of three datasets is used as a metric here. The results showed that even the FSG-Net-N surpassed the other models. By comparing the scores in Table 2, the FSG-Net-N with parameter size of 1.17M outperformed recent studies with an average rank (5.2) across all metrics on the three datasets, compared to SAU-Net’s rank (10.8) with a parameter size of 0.5M, DCASU-Net’s [https://doi.org/10.48550/arxiv.2202.00972] rank (8.0) with a parameter size of 2.6M, ConvU-NeXt’s [HAN2022109512] rank (6.8) with a parameter size of 3.5M, and FR-UNet’s rank (8.2) with a parameter size of 7.4M.
Table 6: Ablation study on the proposed modules. The abbreviations represent: DC – replacement of the residual block with a down-convolution structure; GRM – integration of the GRM block; SA – spatial attention at the bottleneck; DKS – dynamic kernel size evaluation; DS – use of deep supervision.
  DC GRM SA DKS DS Params(M) DRIVE STARE CHASE_DB1
39.09 82.580 81.786 78.944
16.85 82.254 83.323 79.409
18.31 83.115 84.671 79.333
18.32 83.131 84.698 79.889
35.77 83.006 84.200 80.044
18.32 83.229 85.100 81.019
To validate the contribution of each proposed component, we conducted a series of ablation experiments on the DRIVE, STARE, and CHASE_DB1 datasets. Table 6 presents the ablation study conducted on the main modules of FSG-Net, with each module denoted by its corresponding abbreviation. Starting from the R2U-Net [https://doi.org/10.48550/arxiv.1802.06955] baseline, chosen for its strong generalization across medical image segmentation tasks, we incrementally integrated the modules proposed in FSG-Net. As shown in the second row of Table 6, replacing the residual blocks in R2U-Net with our proposed down-convolution (as illustrated in Fig. 3(c)) led to a slight performance drop on the DRIVE dataset, while improving the results on STARE and CHASE_DB1. However, this modification reduced the number of parameters by more than half, indicating a favorable trade-off between efficiency and accuracy. As can be seen in the third row of the Table 6, substituting the standard U-Net-style feature concatenation with the proposed GRM block resulted in notable performance improvements on both DRIVE and STARE. This confirms that our GRM module enhances multi-scale feature aggregation and representation learning. Introducing a lightweight spatial attention mechanism at the bottleneck stage yielded consistent performance gains with only a marginal increase in parameter count. This demonstrates the module’s effectiveness in enhancing contextual understanding without significant overhead. To assess the impact of kernel size, we compared a series of three 3×3 convolutions with a single 7×7 convolution. While the 7×7 configuration exhibited minor improvements in high-resolution scenarios due to its direct abstraction style, it generally led to increased parameter counts and lower performance on other datasets, suggesting limited generalizability. Incorporating deep supervision, as shown in the sixth row of the Table 6, consistently improved performance, supporting its role in guiding the learning process through additional gradient signals during training.
Discussion
Figure 7: Compared structure of feature merging in (a) FSG-Net, and (b) AG-Net.
We previously emphasized the method of feature merging within the network architecture. This is because feature merging allows information from all stages to be mixed and provided as input to the guided filtering module. This approach can be compared to AG-Net, which also gathers features and provides them to the guided filter. Fig. 7(a) illustrates the feature merging in the proposed method, while Fig 7(b) shows the corresponding part of AG-Net. Let us focus on the left path to the guided filtering modules in both cases. As can be seen in Fig 7(a), feature maps from three stages (current, upward, and downward) are concatenated and passed through the proposed convolution block. For AG-Net, only the features from the current and lower stages are merged for the attention-guided filter. In summary, our GRM gets feature information from all stages with a modernized convolution, while the AG-Net gets it from current and lower stages with conventional convolutions. For this reason, we were able to achieve superior performance with one fewer stage than AG-Net.
The next discussion point concerns the experimental datasets. DRIVE, STARE, and Chase_DB1 are representative datasets for retinal vessel segmentation. Models that improve the architecture of layers and attention mechanisms continue to emerge, but the number of images in such datasets is limited. Although this paper focuses on retinal vessel segmentation, it is necessary to apply the proposed model to large databases from other fields to verify its variability and scalability. To this end, we plan to extend FSG-Net to a broader spectrum of medical-imaging tasks—including semi-supervised histological image segmentation [jin2022semi, jin2024inter], skin-lesion diagnosis and segmentation [jin2021cascade], and COVID-19 infection delineation in CT images [jin2021domain]—so as to validate its robustness under limited annotations and pronounced domain shifts.
Finally, we discuss recent approaches to retinal vessel segmentation that differ from ours in terms of network architecture and utilization of attention mechanisms. Unlike our proposed FSG-Net, which utilizes outputs inferred from all layers of a U-Net architecture, Tan et al. [tan2024deep] proposed a W-shaped RCAR-UNet structure that aggregates outputs from the preceding U-Net and passes them forward to the next U-Net module. Meanwhile, Ding et al. [ding2024rcar] incorporated a rough set-based channel-attention mechanism into the U-Net structure to reinforce the long-range dependency of retinal vessels. This differs from our approach, which combines information from all scales and passes it through a trainable guided filter. In future studies, we plan to explore integrating these complementary attention and filtering strategies with our full-scale guided network framework, potentially yielding further improvements in segmentation performance.
Conclusion
In this study, we presented a full-scale representation guided network (FSG-Net) for retinal vessel segmentation that demonstrates competitive performance approaching current SOTA methods on the DRIVE dataset. A modern convolutional block tailored for retinal vessel segmentation was designed, and the guided convolution block effectively leveraged full-scale information from the feature representation network. The proposed guided convolution block is compatible with any U-Net architecture, offering scalability for similar tasks across various domains. However, our method remains purely CNN-based, which might limit its capability compared to recent transformer-integrated U-Net architectures designed for retinal vessel segmentation tasks. Therefore, future research should explore integrating transformer modules into our full-scale feature representation approach. We encourage future research to use the FSG-Net architecture and our experimental settings to ensure reproducibility and scalability in similar tasks.